# A Tidy Data Structure and Visualisations for Multiple Variable Correlations and Other Pairwise Scores

### v1

@@ Abstract
We provide a pipeline for calculating, managing and visualising correlations and other pairwise scores for numerical and categorical data. We present a uniform interface for calculating a plethora of pairwise scores and a new tidy data structure for managing the results. We also provide new visualisations which simultaneously show multiple and/or grouped pairwise scores. The visualisations are far richer than a traditional heatmap of correlation scores, as they help identify relationships with categorical variables, numeric variable pairs with non-linear associations or those which exhibit Simpson’s paradox. These methods are available in our R package bullseye.
@@ 1   Introduction
In this paper we present a new package bullseye (Chinwan and Hurley 2024) which offers a uniform interface for calculating various scores assessing the relationship between variable pairs, both numerical and categorical. We construct a tidy data structure for managing multiple and/or group-wise bivariate scores. We design a new visualisation method which is a variant of the correlation heatmap matrix for showing multiple pairwise or grouped scores. The variant which compares overall to grouped scores is reminiscent of a bullseye, hence the package name. We further design a new linear display, which is useful for higher-dimensional datasets, when interest lies in displaying just interesting pairs of variables. Both displays are optionally interactive.
More general pairwise variable scores known as scagnostics also fit into the framework offered by bullseye. These numeric variable scores were first introduced by Tukey and Tukey (1985) and further developed by Wilkinson, Anand, and Grossman (2005) and seek to identify patterns such as clumpiness and the presence of outliers, for the purpose of identifying scatterplots worthy of detailed exploration.
The outline of the paper is as follows: In Section 2 we present background, first discussing R packages offering visualisations of correlation summaries, and then move to R packages providing calculation of pairwise variable scores beyond the familiar Pearson, Spearman or Kendall’s correlation. In Section 3 we introduce bullseye, describing the new tidy data structure for managing pairwise scores, and the range of functions offered for calculating them. In Section 4 we present the new visualisations offered in bullseye and give some examples. Section 5 describes how bullseye can be used in co-operation with other R related packages. Finally we conclude with a discussion.
@@ 2   Related R packages
D. J. Murdoch and Chow (1996) proposed a display for correlation matrices which uses a matrix layout of ellipses where the parameters of the ellipses are scaled to the correlation values. An implementation is provided in package ellipse (D. Murdoch and Chow 2023). Friendly (2002) expanded on this idea by rendering correlation values as shaded squares, bars, ellipses, or circular pac-man symbols, and these methods are implemented in packages corrplot (Wei and Simko 2021) and corrgram (Wright 2021). The ggcorrplot (Kassambara 2023) and corrr (Kuhn, Jackson, and Cimentada 2024) packages also provided heatmap-type displays, but based on ggplot2 (Wickham 2016) rather than base R.
A number of packages support network displays of correlation values. In the package corrr network displays use line thickness to encode correlation magnitude, with a filtering option to discard low-correlation edges. linkspotter (Samba 2020) offers a variety of association measures in addition to correlation, where the measure used depends on whether the variables are both numerical, categorical or mixed. The results are visualised in a network plot, which may be packaged into an interactive shiny application. The package corrgrapher (Morgen and Biecek 2020) also uses a network plot, and handles mixed type variables by using association measures obtained as transformations of $p$-values obtained from Pearson’s correlation test in the case of two numeric variables, Kruskal’s test for numerical and factor variables, and a chi-squared test for two categorical variables.
The R package correlationfunnel Dancho (2020) offers a novel linear display which assists in feature selection in a regression setting with many predictor variables. All numeric variables including the response are binned. All (now categorical) variables in the resulting dataset are one-hot encoded and Pearson’s correlation calculated with the response categories. The correlations are visualised in a dot-plot display, where predictors are ordered by maximum correlation magnitude, giving a funnel-like shape.
The majority of the listed packages display the result of the standard cor calculation, so the input data structure for the visualisations is a $p\times p$ symmetric matrix of numerical values, where $p$ is the number of variables. corrr uses a $p\times p$ dataframe instead, or alternatively a long format dataframe where each pair of variables form a row, and provides some convenience functions for manipulating these structures. linkspotter similarly uses a long dataframe where pairs of variables form a row, with additional columns for possible multiple correlations. The package correlation (Makowski et al. 2022) also uses a long dataframe, and for grouped correlations each row is associated with a variable pair and group level. It also offers heatmap and network-style visualisations, though not, as far as we can tell, for grouped correlations.
@@
For nominal variable pairs, many classical measures of association are implemented in the R package DescTools (Signorell 2024). One of these is Pearson’s contingency coefficient which uses the ${\chi}^{2}$ value (scaled to the unit interval) from Pearson’s ${\chi}^{2}$ test for independence. Another is the uncertainty coefficient (Theil 1970) which measures the proportion of uncertainty in one variable which is explained by the other. A symmetric version of this coefficient is obtained by taking the mean of the uncertainty coefficients treating each variable as an independent variable once.
@@
DescTools also offers implementations of many association measures for ordinal pairs of variables, including Kendall’s $\tau$ and $\tau_{b}$ and Goodman Kruskal’s gamma, all of which are based on the on the number of concordances and discordances in paired observations (see for example Agresti (2010) for a description). polycor (Fox 2022) implements polychoric correlation (Olsson 1979), which summarizes the association between two ordinal variables by assuming two normally distributed latent variables. It also provides polyserial correlation, a measure of association between an ordinal, numeric pair of variables.
@@
For numeric pairs of variables, the recently proposed measures distance correlation (Székely, Rizzo, and Bakirov 2007) and maximal information coefficient (MIC) (Reshef et al. 2011) are measures of dependence which aim to capture non-linear as well as linear patterns. Distance correlation is calculated as the correlation between two sets of distances, and is available from the R package energy (Rizzo and Szekely 2022). The MIC coefficient is calculated as the mutual information between discretised versions of the two numeric variables, where the binning is chosen to maximise the mutual information. Simon and Tibshirani (2014) simulated pairs of variables with different relationships at varying levels of noise and showed that distance correlation has more statistical power than MIC. In a later study, Kinney and Atwal (2014) proposed additional mutual information-based measures with improved performance. The MIC and other maximal information based non-parametric exploration statistics have been implemented in the R package minerva (Albanese et al. 2012). Another package linkspotter provides a mutual information measure for pairs of variables which may both be numeric, factors or mixed, which is similar to the minerva MIC measure for numeric pairs.
@@
The alternating conditional expectations (ACE) algorithm (Breiman and Friedman 1985) as implemented in acepack (Spector et al. 2016) estimates optimal transformations between response and predictor variables for regression analysis and calculates $R^{2}$ using this optimal transformation. For a numeric pair of variables, this provides another way of summarizing their non-linear association. For categorical variables, ACE transformations calculates the optimal scoring of levels (also known as canonical analysis), and the resulting $R^{2}$ provides an alternative measure of association.
@@
The package scagnostics (Wilkinson and Anand 2022) implements the methods in Wilkinson, Anand, and Grossman (2005) evaluating scatterplot “interestingness”, and provides nine scores called outlying, skewed, clumpy, sparse, striated, convex, skinny, stringy and monotonic.
3   Bullseye calculations
In this section and the following Section 4 we describe the main features of our new package bullseye. Here we present a tidy data structure for managing pairwise scores, which offers far more flexibility than the standard correlation matrix. Then we describe the functions included for calculating individual, multiple and grouped pairwise scores. In the next section we discuss the new visualisations offered and give some examples.
For the examples in this section we use the penguins dataset from package palmerpenguins (Horst, Hill, and Gorman 2020), which lists measurements for penguin species from three islands in the Palmer Archipelago. The dataset has five numerical variables, and three factor variables species, island and sex.
@@ 3   Bullseye calculations
Conventionally, correlation measures are represented as a matrix, following the mathematical definition of a correlation matrix. This matrix representation does not follow tidy data principles and does not extend naturally to the situation of multiple or grouped pairwise measures. For these reasons, some existing packages (for example corrr, linkspotter and correlation) provide tidy dataframe representations.
@@
For our purposes, we design a pairwise structure that contains as columns (i) the two variables, (ii) the score name (iii) the group (iv) the score value and (v) the types of the variables. Starting from a correlation matrix of numeric variables in the penguins dataset, constructed below as peng_cor, the pairwise function constructs the equivalent pairwise structure:
@@
As bullseye is designed for symmetric measures, each variable pair appears once only. By construction the x and y variables are in alphabetical order, that is, x < y. Note there is no row where x = y. Here the correlations are not grouped so group="all". In the pair_type column, we use nn for two numeric variables, ff for two factors and fn for a factor-numeric pair. The above calculation can be done in one step with the bullseye function pair_cor. Notice that this automatically excludes non-numeric variables and handles missing values using pairwise complete observations. In the example below, the optional argument method = "spearman" is specified so Spearman correlation is calculated, rather than the default Pearson.
pair_cor(penguins, method="spearman")
#> # A tibble: 10 x 6
#>   x         y         score    group  value pair_type
#>   <chr>     <chr>     <chr>    <chr>  <dbl> <chr>
#> 1 bill_dep  bill_len  spearman all   -0.222 nn
#> 2 bill_len  flip_len  spearman all    0.673 nn
#> 3 bill_dep  flip_len  spearman all   -0.523 nn
#> 4 body_mass flip_len  spearman all    0.840 nn
#> 5 bill_len  body_mass spearman all    0.584 nn
#> 6 bill_dep  body_mass spearman all   -0.432 nn
#> # i 4 more rows
Calculating pairwise scores
Table 2 has a list of functions provided by bullseye to calculate pairwise variable scores. Each of these functions takes as input a data.frame containing a dataset and handle.na which defaults to TRUE (and uses pairwise complete observations). Each excludes variables which are not of the appropriate type for the score to be calculated. Other arguments may optionally be passed to the underlying calculation function, specified in the “From” column of Table 2. Each of these pair_xx functions produces an output which is of class pairwise, with the score and pair_type columns filled in as appropriate. The function pair_mine will optionally provide multiple measures for each variable pair, while pair_scagnostics defaults to providing all nine scagnostics from scagnostics.
Table 2: Bullseye functions for calculating pairwise corrrelations, other association measures and scagnostics. Tick marks in the columns nn, ff, fn indicate whether the function gives a score for numeric, factor and factor-numeric pairs. Ticks in the Ordinal column indicates method assumes ordinal factors.
@@
We previously saw an example of pair_cor in action. Here is pair_mine, where two methods are specified. The default choice here is just method="MIC".
@@
pairwise_multi Combines outputs of pair_xx functions.
pairwise_by Combines outputs of a pair_ xx function for every level of a grouping variable.
pairwise_scores Combines outputs of pair_ xx functions, where different pair_xx functions may be specified for nn, fn, and ff pairs. A grouping variable may optionally be specified.
@@
For each of the pairwise calculation functions, they can be wrapped using pairwise_by to build a score calculation for each level of a grouping variable. (Of course, grouped scores could be calculated using dplyr (Wickham et al. 2023) machinery, but it is a bit more work.) In the code below, correlations are calculated for each penguin species, in addition to the overall or ungrouped correlation. Supplying argument ungrouped=FALSE suppresses calculation of the ungrouped scores.
@@
The default uses Pearson’s correlation for numeric pairs, and canonical correlation for factor-numeric or factor-factor pairs. In addition, polychoric correlation is used for two ordered factors, but there are no ordered factors in this data. Alternative scores may be specified using the control argument to pairwise_scores. pairwise_scores also has a by argument, and in this way provides pairwise scores for the levels of a grouping variable. Once again, use ungrouped=FALSE to suppress calculation of the ungrouped scores.
pairwise_scores(penguins, by="species")
#> # A tibble: 84 x 6
#>   x         y      score  group   value pair_type
#>   <chr>     <chr>  <chr>  <fct>   <dbl> <chr>
#> 1 island    sex    cancor Adelie 0.0164 ff
#> 2 bill_len  island cancor Adelie 0.0838 fn
#> 3 bill_dep  island cancor Adelie 0.0629 fn
#> 4 flip_len  island cancor Adelie 0.148  fn
#> 5 body_mass island cancor Adelie 0.0208 fn
#> 6 bill_len  sex    cancor Adelie 0.590  fn
@@
4   Bullseye visualisations
In this section we describe the visualisation functions plot_pairwise and plot_pairwise_linear offered by bullseye. Then we present some examples.
Visualisation functions
The first visualisation function plot_pairwise is the default plot method for objects of class pairwise. It provides a traditional matrix layout, generalising the usual heatmap. For each (x,y) pair a circular glyph is drawn, divided into equal-sized wedges if the (x,y) pair appears multiple times. Each wedge has a fill level according to the associated value. To distinguish rows representing the complete dataset from those representing subsets, the circular glyph is divided into an inner circle (the bullseye) and an outer doughnut. Wedges in the inner circle represent the complete dataset (identified as rows with group = "all"), and segments of the doughnut represent dataset subsets. In an earlier variant of this display (see Chinwan (2024) Figure 3.4) , we used lollipops to show the scores in each matrix cell, but we find the bullseye version more impactful, though admittedly this assessment was not the subject of an experiment.
Seriation is used to order variable pairs to help draw attention to high-score pairs. In the case of a single score for each (x,y) pair, the usual way of ordering variables employs hierarchical clustering to place high-score (typically correlation) pairs nearby. This has the effect of placing high-score pairs of variables near the diagonal of the matrix layout. In plot_pairwise we use a variant of this scheme, which additionally aims to give high scoring pairs extra prominence by placing them in the top-left corner, using the Earle and Hurley (2015) lazy path length algorithm implemented in package DendSer (dendser?). For multiple scores, for each (x,y) pair we use the maximum of the absolute values, or alternatively the maximum difference.
The second visualisation function plot_pairwise_linear is also available using the plot method for objects of class pairwise, but with type="linear". Each row of the display shows all the values for a given (x,y) pair, using either a tile or point geom. This display is useful for datasets with lots of variables, where the pairwise structure can optionally be filtered to show interesting (x,y) pairs only, or in the case of a regression problem, pairs where a specified response variable is present. Once again, seriation is used to order variable pairs to help draw attention to high-score pairs. Seriation for this linear display is more straightforward than that for the matrix display. In plot_pairwise_linear variable pairs are ordered by their maximum absolute value, or alternatively the maximum difference.
Both visualisations functions produce ggplot2 (ggplot2) objects (Wickham 2016). Alternatively, they have an interactive option, in which case package ggiraph (Gohel and Skintzos 2024) is used and tooltips provide more information on geom objects.
Visualisation examples
@@ 4   Bullseye visualisations
plot(pairwise_scores(penguins))
@@
Figure 1: Pairwise correlations for numerical variables and canonical correlation for factors and mixed factor-numeric pairs. Notice the categorical variable species is strongly associated with the penguin dimension variables, and with the categorical variable island.
@@
The above code calculates pairwise_scores for the penguins data, Figure 1 shows the resulting display. As the variables species, sex and island are factors, scores involving any of these variables is a canonical correlation, while Pearson’s correlation is shown for numerical pairs. As the species variable is strongly associated with the measurement variables, it is clear that the size varies across the penguin species. The species variable is also strongly associated with island, telling us that the species distribution varies across the islands. Not surprisingly, year has little association with the other variables.

### v2

@@ Abstract
We provide a pipeline for calculating, managing and visualising correlations and other pairwise association scores for numerical and categorical data. We present a uniform interface for calculating a plethora of pairwise scores and propose a tidy data structure for organising the results. We also provide new visualisations which simultaneously show multiple and/or grouped pairwise scores. The visualisations are far richer than a traditional heatmap of correlation scores, as they help identify relationships with categorical variables, numeric variable pairs with non-linear associations or those which exhibit Simpson’s paradox. These methods are available in our R package bullseye.
@@ 1  Introduction
In this paper we present a new package bullseye (bullseye) which offers a uniform interface for calculating various scores assessing the relationship between variable pairs, both numerical and categorical. We construct a tidy data structure for managing multiple and/or group-wise bivariate scores. We design a new visualisation method which is a variant of the correlation heatmap matrix for showing multiple pairwise or grouped scores. The variant which compares overall to grouped scores is reminiscent of a bullseye, hence the package name. We further design a new linear display, which is useful for higher-dimensional datasets, when interest lies in displaying just interesting pairs of variables. Both displays are optionally interactive.
More general pairwise variable scores known as scagnostics also fit into the framework offered by bullseye. These numeric variable scores were first introduced by tukey1985computer and further developed by wilkinson2005graph and seek to identify patterns such as clumpiness and the presence of outliers, for the purpose of identifying scatterplots worthy of detailed exploration.
The outline of the paper is as follows: In Section 2 we present background, first discussing R packages offering visualisations of correlation summaries, and then move to R packages providing calculation of pairwise variable scores beyond the familiar Pearson, Spearman or Kendall’s correlation. In Section 3 we describe and motivate the choices made in designing a tidy data structure for representing association summaries, and the new displays for their visualisation. In Section 4 we describe our package bullseye, describing the range of functions for calculating and visualising pairwise associations and give some examples. Section 5 describes how bullseye can be used in co-operation with other R related packages. Finally we conclude with a discussion.
@@ 2  Related R packages
murdoch1996 proposed a display for correlation matrices which uses a matrix layout of ellipses where the parameters of the ellipses are scaled to the correlation values. An implementation is provided in package ellipse (ellipse). friendly2002corrgrams expanded on this idea by rendering correlation values as shaded squares, bars, ellipses, or circular pac-man symbols, and these methods are implemented in packages corrplot (corrplot) and corrgram (corrgram). The ggcorrplot (ggcorrplot) and corrr (corrr) packages also provided heatmap-type displays, but based on ggplot2 (ggplot2) rather than base R. FriendlySigal (in their Figure 2) present displays of grouped correlations as overlaid ellipses. This is implemented in the function covEllipse of package heplots (heplots).
A number of packages support network displays of correlation values. In the package corrr network displays use line thickness to encode correlation magnitude, with a filtering option to discard low-correlation edges. linkspotter (linkspotter) offers a variety of association measures in addition to correlation, where the measure used depends on whether the variables are both numerical, categorical or mixed. The results are visualised in a network plot, which may be packaged into an interactive shiny application. The package corrgrapher (corrgrapher) also uses a network plot, and handles mixed type variables by using association measures obtained as transformations of $p$-values obtained from Pearson’s correlation test in the case of two numeric variables, Kruskal’s test for numerical and factor variables, and a chi-squared test for two categorical variables.
The R package correlationfunnel correlationfunnel offers a novel linear display which assists in feature selection in a regression setting with many predictor variables. All numeric variables including the response are binned. All (now categorical) variables in the resulting dataset are one-hot encoded and Pearson’s correlation calculated with the response categories. The correlations are visualised in a dot-plot display, where predictors are ordered by maximum correlation magnitude, giving a funnel-like shape.
The majority of the listed packages display the result of the standard cor calculation, so the input data structure for the visualisations is a $p\times p$ symmetric matrix of numerical values, where $p$ is the number of variables. corrr uses a $p\times p$ dataframe instead, or alternatively a long format dataframe where each pair of variables form a row, and provides some convenience functions for manipulating these structures. linkspotter similarly uses a long dataframe where pairs of variables form a row, with additional columns for possible multiple correlations. The package correlation (correlationPackage) also uses a long dataframe, and for grouped correlations each row is associated with a variable pair and group level. It provides heatmap and network-style visualisations, though not, as far as we can tell, for grouped correlations.
@@
For nominal variable pairs, many classical measures of association are implemented in the R package DescTools (DescTools). One of these is Pearson’s contingency coefficient which uses the ${\chi}^{2}$ value (scaled to the unit interval) from Pearson’s ${\chi}^{2}$ test for independence. Another is the uncertainty coefficient (theil1970estimation) which measures the proportion of uncertainty in one variable which is explained by the other. A symmetric version of this coefficient is obtained by taking the mean of the uncertainty coefficients treating each variable as an independent variable once.
@@
DescTools also offers implementations of many association measures for ordinal pairs of variables, including Kendall’s $\tau$ and $\tau_{b}$ and Goodman Kruskal’s gamma, all of which are based on the on the number of concordances and discordances in paired observations (see for example agresti2010analysis for a description). polycor (polycor) implements polychoric correlation (olsson1979maximum), which summarizes the association between two ordinal variables by assuming two normally distributed latent variables. It also provides polyserial correlation, a measure of association between an ordinal, numeric pair of variables.
@@
For numeric pairs of variables, the recently proposed measures distance correlation (szekely2007measuring) and maximal information coefficient (MIC) (reshef2011detecting) are measures of dependence which aim to capture non-linear as well as linear patterns. Distance correlation is calculated as the correlation between two sets of distances, and is available from the R package energy (energy). The MIC coefficient is calculated as the mutual information between discretised versions of the two numeric variables, where the binning is chosen to maximise the mutual information. commentSimonTibshirani simulated pairs of variables with different relationships at varying levels of noise and showed that distance correlation has more statistical power than MIC. In a later study, kinney proposed additional mutual information-based measures with improved performance. The MIC and other maximal information based non-parametric exploration statistics have been implemented in the R package minerva (minerva). Another package linkspotter provides a mutual information measure for pairs of variables which may both be numeric, factors or mixed, which is similar to the minerva MIC measure for numeric pairs.
@@
The alternating conditional expectations (ACE) algorithm (breiman1985estimating) as implemented in acepack (acepack) estimates optimal transformations between response and predictor variables for regression analysis and calculates $R^{2}$ using this optimal transformation. For a numeric pair of variables, this provides another way of summarizing their non-linear association. For categorical variables, ACE transformations calculates the optimal scoring of levels (also known as canonical analysis), and the resulting $R^{2}$ provides an alternative measure of association.
@@
The package scagnostics (scagnostics) implements the methods in wilkinson2005graph evaluating scatterplot “interestingness”, and provides nine scores called outlying, skewed, clumpy, sparse, striated, convex, skinny, stringy and monotonic.
3  Bullseye design
In this section we present design choices made in constructing the new package bullseye.
We firstly describe a tidy data structure for managing pairwise scores, which offers far more flexibility than the standard correlation matrix. While we anticipate most users of the package will be more interested in the visualisations provided, understanding this structure will be of interest to those who wish to use bullseye calculations with visualisations from other packages, or conversely those who wish to use association measures from other packages with bullseye visualisaions.
Then we describe and motivate our new visualisations.
@@ 3  Bullseye design
Conventionally, correlation measures are represented as a matrix, following the mathematical definition of a correlation matrix. This matrix representation does not follow tidy data principles as described in (wickhamjss). Furthermore, the matrix arrangement does not extend naturally to the situation of multiple or grouped pairwise measures. For these reasons, some existing packages (for example corrr, linkspotter and correlation) provide tidy dataframe representations.
Organising data according to tidy data principles means variables are columns and observations are rows. For example, a correlation matrix could be converted to a tidy format by having two variables for the row and column label of the matrix, and a third variable for the correlation value.
For our purposes, we use a tibble structure that contains as variables:
  1. (i)
    the two dataset variables x and y,
  2. (ii)
    score: the score name identifying the summary measure used for the two variables,
  3. (iii)
    group: identifying the subset of the data used for calculating the summary and
  4. (iv)
    value: a numeric value recording the value of the summary measure.
As the pairwise measures we consider are symmetric, each variable pair appears once only. By construction the x and y variables are in alphabetical order, that is, x < y. Note there is no row where x = y. An additional column pair_type is included for information purposes. Here we use nn for two numeric variables, ff for two factors and fn for a factor-numeric pair, or NA if unknown. The variables x, y, score, group uniquely identify each row, that is, the pairwise tibble has at most one row for each combination of x, y, score and group. This requirement will be checked prior to using bullseye visualisaions.
Design of visualisations
Ideally, we would design a visualisation where each of x, y, score, group and value columns of the pairwise tibble are mapped to a visual channel, that is, some aspect of the plot. cf13 identified over 30 visual channels, most of which have not been subjected to detailed comparisons in the literature. cleveland84 proposed a ranking of 10 channels from most accurate (position along a common scale) to least accurate (colour saturation), and conducted experiments based on two-value ratio judgements to verify the ranking, though the experiments excluded colour. borgo13 reported on perceptual studies where the “pop-out” effects of colour are far greater than those of other channels such as size, shape and orientation. More recently, coleman22 conducted extensive experiments on six visual channels, including luminance. For the task of reproducing up to eight values, they found that the rankings of cleveland84 did not hold, and that performance is vastly more affected by the number of items to be compared than the channel used. Of the six visual channels explored, luminance was ranked mid table across the various settings. They concluded that much more work needs to be done to rank visual channels across a range of scenarios.
Matrix display
A natural starting point for the visualisation of the pairwise tibble would be a generalisation of the familiar heatmap or glyph representation of a matrix, such as displays provided by many of the packages listed in Table 1. Then the x and y variables are mapped to layout positions, and the other variables score, group and value are mapped to a glyph. borgo13 and Ward2008 both give useful surveys of glyph-based design.
In designing the visualisation, note that the pairwise tibble could have:
  1. a.
    one measure and value for each x, y pair,
  2. b.
    multiple measures, groups and values for each x,y pair, or
  3. c.
    multiple measures, groups and values for each x,y pair, where there is a measure for group = "all" and others for subsets.
The first case (a) is the simplest, which is addressed by the matrix displays of packages listed in Table 1. The most popular display is a matrix of filled squares or circles, using a diverging colour scale, where colour hue represents positive or negative values, and other colour aspects (chroma and luminance) represent the magnitude. This is our choice of display in bullseye for this setting (see Figure 1). Generally, the analyst uses these matrix displays to easily identify pairs of variables with strong, weak or negative associations, rather than for accurate reading of precise association values, so Cleveland and McGill’s hierarchy of visual channels is less of a concern.
For case (b), we decided to split the filled circle into sectors of equal size, one for each value, where each sector is assigned a colour representing a value. See Figure 2 for an example. With this visualisation, the pairs of variables with strong, weak or negative associations, and pairs of variables with large differences across the different values “pop-out” strongly.
For the special case of (c) where some measures represent all the data and others represent subsets, we divide the circle into an inner section whose sectors show ungrouped score values for group = "all", and the outer doughnut whose sectors shows the score values for other groups. As before, sectors are assigned a colour representing the value sign and magnitude. This glyph design represents the hierarchical nature of the values when some summarise the entire dataset and others summarise the data which has been split into subsets. An example is given in Figure 3. An interesting pattern in this context is where the direction of the association is reversed when data is split into groups (that is, Simpson’s paradox), and in our visualisation, this pattern is easily identified by distinct hues for the inner and outer sectors.
Ward2008 also discusses the role of ordering, or seriation, in constructing more informative glyph displays. For our matrix displays, there are two possible uses of seriation (i) ordering the (x,y) pairs and (ii) ordering the slices within each glyph. At the moment we do not implement a strategy for (ii). For (i), we use seriation to order variable pairs to help draw attention to interesting glyphs. In the case of a single score for each (x,y) pair, the usual way of ordering variables employs hierarchical clustering to place high-score (typically correlation) pairs nearby. This has the effect of placing high-score pairs of variables near the diagonal of the matrix layout. We use a variant of this scheme, which additionally aims to give high scoring pairs extra prominence by placing them in the top-left corner, using the dendserPaper lazy path length algorithm implemented in package DendSer (DendSer). For cases (b) and (c) with multiple scores, interesting pairs of variables will likely be those with one or more high scores, or alternatively those high variation across the scores. In these settings our seriation procedure arranges the variable pairs based on the the maximum of the absolute values, or alternatively the maximum difference of those values.
In a previous version of this work (chinwan) (see for example Figure 3.4 and 3.6) we experimented with representing values by lines using position along a common scale, the most accurate representation according to cleveland84. However, the display did not make good use of available space and in our opinion did not have the same “pop-out” effect as our current designs, though this assessment has not been subjected to experimentation.
Note that in the matrix layout with the divided circle glyph the group and score variables are mapped to sector position, making it difficult for the user to see which sector is which. This is a problem common to multivariate glyphs in general, for example stars and Chernoff faces, and as discussed in Ward2008, can lead to bias in glyph mappings. In our setting, we have opted to add interactive tool tips to the displays so that the group and or score as well as the precise value associated with each sector is easily identified.
Linear display
The matrix display described above is suitable when the number of variables is not too large. As the number of variables increases, the pairwise tibble could be filtered to show interesting (x,y) pairs only, or in the case of a regression problem, pairs where a specified response variable is present. For this setting we use a dotplot display, where (x,y) pairs are mapped to rows, value is mapped to position, and the group/score categories are mapped to point colour. We also offer a heatmap variant where value is mapped to tile fill. See Figures 4 and 5 for examples.
Once again, seriation is used to order variable pairs to help draw attention to high-score pairs. Here variable pairs are ordered by their maximum absolute value, or alternatively the maximum difference.
4  Bullseye functionality
In this section we describe the main functions of our new package bullseye. First we describe the functions included for calculating individual, multiple and grouped pairwise scores. Then we present visualisation functions and give some examples.
We demonstrate the functions using the penguins dataset from package palmerpenguins (palmerpenguins), which lists measurements for penguin species from three islands in the Palmer Archipelago. The dataset has five numerical variables, and three factor variables species, island and sex.
Calculating pairwise scores
@@ 4  Bullseye functionality
Starting from a correlation matrix of numeric variables in the penguins dataset, constructed below as peng_cor, the pairwise function constructs the equivalent pairwise tibble:
@@
We also provide utility functions to calculate various pairwise variable scores. These are listed in Table 2. Each of these functions takes as input a data.frame containing a dataset and handle.na which defaults to TRUE (and uses pairwise complete observations). Each excludes variables which are not of the appropriate type for the score to be calculated. Other arguments may optionally be passed to the underlying calculation function, specified in the “From” column of Table 2. Each of these pair_* functions produces an output which is a tibble of class pairwise, with the score and pair_type columns filled in as appropriate. The function pair_mine will optionally provide multiple measures for each variable pair, while pair_scagnostics defaults to providing all nine scagnostics from scagnostics.
Table 2: Bullseye functions for calculating pairwise correlations, other association measures and scagnostics. Tick marks in the columns $n$, $f$, fn indicate whether the function gives a score for numeric, factor and factor-numeric pairs. Ticks in the Ordinal column indicates method assumes ordinal factors.
@@
The calculation in the previous code snippet can be done in one step with the bullseye function pair_cor. Notice that this automatically excludes non-numeric variables and handles missing values using pairwise complete observations. In the example below, the optional argument method = "spearman" is specified so Spearman correlation is calculated, rather than the default Pearson.
pair_cor(penguins, method="spearman")
#> # A tibble: 10 x 6
#>   x         y         score    group  value pair_type
#>   <chr>     <chr>     <chr>    <chr>  <dbl> <chr>
#> 1 bill_dep  bill_len  spearman all   -0.222 nn
#> 2 bill_len  flip_len  spearman all    0.673 nn
#> 3 bill_dep  flip_len  spearman all   -0.523 nn
#> 4 body_mass flip_len  spearman all    0.840 nn
#> 5 bill_len  body_mass spearman all    0.584 nn
#> 6 bill_dep  body_mass spearman all   -0.432 nn
#> # i 4 more rows
Here is pair_mine, where two methods are specified. The default choice here is just method="MIC".
@@
pairwise_multi Combines outputs of pair_* functions.
pairwise_by Combines outputs of a pair_* function for every level of a grouping variable.
pairwise_scores Combines outputs of pair_* functions, where different pair_* functions may be specified for nn, fn, and ff pairs. A grouping variable may optionally be specified.
@@
For each of the pairwise calculation functions, they can be wrapped using pairwise_by to build a score calculation for each level of a grouping variable. (Of course, grouped scores could be calculated using dplyr (dplyr) machinery, but it is a bit more work.) In the code below, correlations are calculated for each penguin species, in addition to the overall or ungrouped correlation. Supplying argument ungrouped=FALSE suppresses calculation of the ungrouped scores.
@@
The default uses Pearson’s correlation for numeric pairs, and canonical correlation for factor-numeric or factor-factor pairs. In addition, polychoric correlation is used for two ordered factors, but there are no ordered factors in this data. Alternative scores may be specified using the control argument to pairwise_scores. pairwise_scores also has a by argument, and in this way provides pairwise scores for the levels of a grouping variable. Once again, use ungrouped=FALSE to suppress calculation of the ungrouped scores. The two functions pairwise_scores and pairwise_by that calculate grouped scores have an additional argument add.nobs, which when TRUE the returned pairwise tibble has an additional column n containing the number of observations used for each score.
pairwise_scores(penguins, by="species", add.nobs=TRUE)
#> # A tibble: 84 x 7
#>   x         y      score  group   value pair_type     n
#>   <chr>     <chr>  <chr>  <fct>   <dbl> <chr>     <int>
#> 1 island    sex    cancor Adelie 0.0164 ff          146
#> 2 bill_len  island cancor Adelie 0.0838 fn          151
#> 3 bill_dep  island cancor Adelie 0.0629 fn          151
#> 4 flip_len  island cancor Adelie 0.148  fn          151
#> 5 body_mass island cancor Adelie 0.0208 fn          151
#> 6 bill_len  sex    cancor Adelie 0.590  fn          146
@@
Visualising pairwise scores
In this section we describe the visualisation functions offered by bullseye. Then we present some examples.
The first visualisation function plot_pairwise is the default plot method for objects of class pairwise. It provides a traditional matrix layout, generalising the usual heatmap to the setting where there are multiple association values for each (x,y) pair of variables, covering each of the settings (a), (b), (c) listed in Section 3 The second visualisation function plot_pairwise_linear is also available using the plot method for objects of class pairwise, but with type="linear". Each row of the display shows all the values for a given (x,y) pair, using either a tile or point geom.
Both visualisations functions produce ggplot2 (ggplot2) objects (ggplot2). Alternatively, they have an interactive option, in which case package ggiraph (ggiraph) is used and tooltips provide more information on geom objects.
@@
pairwise_scores(penguins) |>
  plot_pairwise()
@@
Figure 1: Pairwise correlations for numerical variables and canonical correlation for factors and mixed factor-numeric pairs from the penguin dataset. Notice the categorical variable species is strongly associated with the penguin dimension variables, and with the categorical variable island.
@@
The above code calculates and plots pairwise_scores for the penguins data, Figure 1 shows the result. As the variables species, sex and island are factors, scores involving any of these variables is a canonical correlation, while Pearson’s correlation is shown for numerical pairs. As the species variable is strongly associated with the measurement variables, it is clear that the size varies across the penguin species. The species variable is also strongly associated with island, telling us that the species distribution varies across the islands. Not surprisingly, year has little association with the other variables.
We now take a more detailed look at the associations between numerical variables, comparing Pearson, Spearman and distance correlation. As we are interested in comparing magnitudes, we take absolute values for the signed measures Pearson and Spearman, see Figure 2. As the three association measures give similar values, there is very little evidence of non-linear association between these four variables.
We have chosen to use (hue, chroma, luminance) HCL-based palettes provided by colorspace (colorspace), for their good perceptual properties. The default palette is "Blue-Red 3" (see Figure 1) which uses the two hues blue and red for negative and positive values, zero values are shown in white, with value magnitude represented by varying chroma and luminance. Alternatively any sequential or diverging palette from colorspace may be specified to the pal argument of plot_pairwise, as in the following code snippet, which uses pal="Purples 2".
penguins |>
  select(bill_len:body_mass) |>
  pairwise_multi(c("pair_cor", "pair_spearman", "pair_dcor")) |>
  mutate(value=abs(value)) |>
  plot_pairwise(pal="Purples 2")
Figure 2: Pearson, Spearman and distance correlation for a subset of variables from the penguin dataset. The three measures are similar in magnitude for all pairs of variables. Tooltips show the score names and values in the html version.